OpenStudy (anonymous):

x-2/y=2 solve for y i got: x-2/2=y am i right?

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

If the equation is \[\Large \frac{x-2}{y}=2\], then the answer is \[\Large y = \frac{x-2}{2}\]

OpenStudy (anonymous):

sorry that x-2/2 is supposed to be x-2/z

jimthompson5910 (jim_thompson5910):

oh

OpenStudy (anonymous):

yes so same answer except for the 2 is a z

jimthompson5910 (jim_thompson5910):

problem is \[\Large \frac{x-2}{y}=z\] ???